OpenStudy (anonymous):

URGENT PLEASE HELP I WILL GIVE MEDALS AND ANYTHING YOU WANT PLEASE VERY IMPORTANT HARD MATH QUESTION!!!!!

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (superflytyguy):

What's up

OpenStudy (anonymous):

Ok. In the early days of the civil war, a soldier had to carry the following equipment while marching: ten pounds of gun equipment including eighty rounds of ball cartridge, one pound of powder and five pounds of lead, knapsack, haversack, three-pint canteen, all full with 3 days rations, rubber blanket, woolen blanket, shelter tent, full winter clothing, tin cup, tin plate, knife, fork, spoon, stationery, photographs, journal, Bible, tobacco, pipes, comb and brush, shaving tools, sewing kit, toothbrush, soap and whatever other gear did not hang from hooks on the soldiers' belts. Throughout the war, these men marched long and hard, sometimes up to 30 or 40 miles a day. If a soldier marched 30 miles in 18 hours, what was his average rate of speed? The formula "d=rt" gives the distance D traveled in time T at a rate R. Use the 5 step organizer to solve the problem above. Answer all parts of the organizer in the box below. Label your answer with Part A, Part B, Part C, Part D, and Part E. A. Identify: (3 points) Read the problem carefully. You may need to read it more than once. What are you being asked to figure out? B. Strategize: (3 points)Use questions like these in planning the solution: Write all of the given information that will help you solve this problem. Are unnecessary facts given? If so, what are they? C. Set up: (3 points) Write the formula that you will use to solve this problem. D. Solve: (3 points) Substitute given values into the formula. Carry out the operations carefully. Show all of your work. E. Check: (3 points) Check your results with the facts given in the problem. State your answer in a complete sentence. Does your answer make sense to you? If not, re-work the problem.

OpenStudy (anonymous):

Is this a fuking project?

OpenStudy (anonymous):

Um no @theslasher00

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

jk jk

OpenStudy (anonymous):

I'm in 7th grade

OpenStudy (anonymous):

@AlexandervonHumboldt2 and @SuperflyTyGuy pleaase help

OpenStudy (anonymous):

#Toeasy but you didn't ask for my help so PEACE OUT ;)

OpenStudy (superflytyguy):

r= d/t, so the rate will be 30/18. this is the same as 5/3, or 1.66666666666miles/hour @Rocky_

OpenStudy (anonymous):

Thanks? Lol @SuperflyTyGuy

OpenStudy (superflytyguy):

Yep

OpenStudy (anonymous):

is that for a b c d or e @SuperflyTyGuy

OpenStudy (superflytyguy):

D

OpenStudy (superflytyguy):

C and D really. I wrote down the method they provided d = rt, switched it around, and then solved it

OpenStudy (superflytyguy):

switched it around r = d/t

OpenStudy (anonymous):

What about a and b? @SuperflyTyGuy

OpenStudy (superflytyguy):

A. What are you being asked to figure out? How fast did he walk. B. Write all of the given information that will help you solve this problem. A soldier marched 30 miles in 18 hours The info about what he was carrying was unnecessary and just to throw you off.

OpenStudy (anonymous):

Thank you so much @SuperflyTyGuy

OpenStudy (superflytyguy):

No problem @Rocky_ :)

OpenStudy (anonymous):

I'm sorry one more thing what is the formula that you figured this out like 1+2=3 @SuperflyTyGuy

OpenStudy (superflytyguy):

r= d/t Rate = Distance divided by time

OpenStudy (anonymous):

Thank you so much